OpenStudy (anonymous):

A car travels at an average speed of 45 miles per hour for a distance of 90 miles. How many minutes did the car travel? A. 45 minutes B. 60 minutes C. 120 minutes D. 135 minutes

OpenStudy (anonymous):

90/45 = 2 hours x 60 = 120 minutes

OpenStudy (anonymous):

@amistre64 Am I right?

OpenStudy (amistre64):

2 hours is 90 miles, i like it

geerky42 (geerky42):

Use the formula: \(v = \dfrac{d}{t}\), where v is speed, d is distance traveled, and t is time. We were already given speed and distance, so we need to find time. \(45mph = \dfrac{90\space miles}{t}\) or just \(45 = \dfrac{90}{t}\) Can you find t?